# Incorporating Procedural Fairness in Flag Submissions on Social Media Platforms

### v1

@@ 5.1.4.   Moderator Type
We examined how different moderator types affect user perceptions. Table 8 shows the mean values of perceived consistency, transparency, and voice in participant groups that encountered the different conditions for moderator types: no information, human, and bot. We then assessed via independent samples t-tests how information (or lack of information) about the moderator responsible for flag review affects users’ fairness perceptions.
Table 8. Mean Values of Perceived Consistency, Transparency, and Voice across Different Moderator Types.
Fairness aspect Moderator type Mean SD SE
Consistency Not available or Human 5.54 1.45 .03
Bot 5.55 1.45 .05
Transparency Not available 5.32 1.38 .04
Present (Human or Bot) 5.30 1.38 .03
Voice Not available or Bot 5.32 1.55 .04
Human 5.26 1.55 .05
  • Total N = 2,936 (Not available = 988, Human = 972, Bot =976)
Perceived Consistency. We first tested whether flagging scenarios depicting a bot moderator were perceived as more consistent than those depicting a human moderator or offering no information about the moderator. The mean value of perceived consistency in the condition where the moderator’s identity is known to be a bot (M = 5.55) is slightly higher than in the condition where the moderator’s identity is either unknown or identified as human (M = 5.54). However, our t-test showed that this difference in perceived consistency between these participants is not significant (t(2934) = 0.27, p = .39).
Perceived Transparency. We next measured whether any information about the moderator’s identity (either bot or human) affects perceived transparency. The mean value of perceived transparency in the condition where the moderator’s identity is not available (M = 5.32) slightly exceeds that in the condition where the moderator’s identity is present (M = 5.30). Yet, our t-test result showed that this difference in perceived transparency is not significant (t(2934) = 0.35, p = .37).
Perceived Voice. Finally, we evaluated whether indications of a human moderator affect perceptions of voice. The mean value of perceived transparency in the condition where the moderator’s identity is either not available or bot (M = 5.32) is slightly higher than in the condition where the moderator’s identity is human (M = 5.26). Still, our t-test showed that this difference in perceived voice is not significant (t(2934) = 1.11, p = .13).
@@
Overall, our analyses found that different moderator identities do not significantly impact perceived consistency, transparency, and voice. Therefore, we reject the hypotheses H4-1, H4-2, and H4-3.
5.1.5.   Additional Analyses: Interaction Effects and Cognitive Burden
@@
Interaction Effects. In addition to testing our proposed hypotheses, we built a General Linear Model (GLM) to examine how different combinations of flagging components affect perceived fairness. Using this model, we explored the interaction effects among the four flagging components on perceived consistency, transparency, and voice. Table 13 in Appendix B shows the results of this analysis. We estimated the statistical significance of these results following Bonferroni correction ($\alpha$05/11) and concluded that the interaction between classification schemes and the availability of a text box significantly affected perceived voice (F(2, 2936) = 8.29, p ¡ .001). All other interactions did not significantly impact any of the three aspects of procedural fairness.
Analyzing this interaction in more detail, our pairwise comparison results indicate that for participants in the condition without a text box, providing either a simple (MD = 0.34, SE = 0.10, p ¡ .001) or detailed (MD = 0.35, SE = 0.10, p ¡ .001) classification scheme elicited significantly higher perceptions of voice than providing no classification scheme (Table 14, Appendix B). However, these participants showed no significant differences in their voice perceptions between the simple and detailed classification scheme conditions (MD = .01, SE = .10, p = 1.00). Further, for participants in the condition with a text box, differences in classification schemes did not significantly impact perceived voice.
Cognitive Effort to Submit Flags. Given that usability is a crucial factor in designing interactive processes like flagging, we examined how different ways of designing each component of flagging mechanisms differentially impacted the cognitive burden on our participants and potentially influenced flag usability.
Components SS df MS f/t p
Classification 8.14 2 4.07 1.31 .27
Guidelines 17.63 2 8.81 2.83 .06
Text box - 2933.87 - -3.03 .01
Moderator 4.01 2 2.00 0.64 .53
Components SS df MS f/t p
Classification .91 2 .46 .19 .83
Guidelines .12 2 .06 .03 .98
Text box - 2930.17 - -1.07 .14
Moderator 8.40 2 4.20 1.77 .17
First, we conducted three separate ANOVA tests to evaluate whether different choices of classification schemes, guidelines, and moderator types impacted participant-reported cognitive burden. Additionally, we performed an independent samples t-test to assess the effect of the text box component on the cognitive burden. Table 10 presents the results of these tests, which revealed that participants in the condition with a text box (M = 3.86, SD = 1.73, EM = .05) experienced a significantly higher cognitive burden compared to those without such a box (M = 3.66, SD = 1.80, EM = .05). No other effects approached statistical significance.
Similarly, we conducted three separate ANOVA tests to evaluate whether different choices of classification schemes, guidelines, and moderator types impacted participants’ self-reported likelihood of using the flagging mechanism in the future. We also performed an independent samples t-test to assess the effect of the text box component on users’ likelihood to flag again. Table 10 presents the results of these tests, showing that variations in how each flagging component is presented (or not presented) did not significantly affect the future use of the flagging mechanism.
In sum, our analysis suggests that offering a text box significantly affects the perceived cognitive burden of the flagging mechanism. However, it does not significantly affect the likelihood of future use of flags. Further, offering additional information about classification schemes, posting guidelines, and moderation types does not compromise the usability of flags.
@@ 5.2. Qualitative Findings
Our inductive analysis identified 5 themes through our combining and distilling of 26 codes that together provide a nuanced understanding of user perspectives on enhancing fairness in flagging mechanisms. Table 11 summarizes the main themes we derived from this analysis along with the frequency of each theme.
@@
Demanding outcome notifications with timely review 362 (29%)
@@ 5.2.1.   Desired Attributes of Flag Reviewers
Participants frequently detailed their preferences for moderators who review the flagged content. Some participants (N = 61) mentioned that they prefer a human reviewer who can understand the context and nuance of the flagged post. Further, a few (N = 9) expressed their interest in communicating with moderators about the flagging process and the results. This suggests that users need thorough explanations for flagging outcomes through detailed conversations, revealing a preference for human interaction. For instance, participant P312 suggested:
@@
Additionally, 17 participants suggested that platforms implement collaborative decision-making by involving multiple moderators for flagged posts. Some participants specified their suggestions for a certain combination or number of moderators (often three) or a mix of human and bot moderators. They believed that such cooperative decision-making in the flagging process would increase fairness and reduce biases. For example, P405 wrote:
@@
Further, to reduce a biased moderation review process, 27 participants emphasized the desired qualifications and characteristics of individual moderators. This includes moderator’s expertise in the subject matter of reviewed content, their understanding of ethical standards regarding rule violations, and their political neutrality.
@@ 5.2.2.   Needing Support for Greater Expressivity
In addition to appreciating the different components they encountered within the reporting mechanism (e.g., posting guidelines, text box), many participants proposed changes to how flagging systems could be better designed. Among these, the most frequently suggested feature was the inclusion of multiple choices for flagging reasons (N = 90), which would enable flaggers to select several rule violation categories within the classification scheme. Our survey was designed to include some examples of rule violations involving multiple infractions, making it challenging for participants to select only one rule violation classification category when the classification scheme was provided. This restriction created a dilemma for some respondents regarding which category they should select, often inducing them to consider how the ability to select multiple options for categorizing rule violations would shape the flagging experience. P343 felt that this feature would help prioritize the flagging of severely inappropriate posts:
For example, I could only choose hate speech, but if it was also considered harassment and could fall under multiple categories, it could be removed or reviewed faster.
Participant P158 similarly advocated for allowing the selection of multiple categories, anticipating that it would induce changes at the flag review stage:
@@
Further, 45 respondents also mentioned preferences for a broader classification scheme with more options for rule violation selections. For instance, P485 wrote:
Implementing a wider range of specific categories for flagging posts.
These responses reflect participants’ desire to classify and elaborate on the types of rule violations they encounter when making their reporting decisions. Other suggestions on how to achieve this included an ability to rate the severity of posts (N = 15) or highlight specific sections of posts that violate rules (N = 14). For instance, P708 noted:
@@
In a related context, some respondents (N = 21) emphasized the importance of discussing toxic posts in a forum where they could share their thoughts on the regulation of those posts. This would facilitate a peer review process, allowing users to specify their reasons for flagging a post and understand others’ perspectives. Participants believed that involving diverse perspectives prior to the flag submission could lead to more accurate and reliable decisions in the post-review stage about whether to remove the flagged content. For instance, P705 commented:
Possibly creating a forum where some users can discuss their objections about a concerning post.
Some participants desired the ability to review similar previously flagged posts and their post-review outcomes before submitting a flag. They felt that this feature, along with the feedback from other users, could ensure that their flags are sufficiently justified for removal by allowing them to more clearly indicate the extent to which flagged content violates rules. Participants felt that their use of these features would bolster the credibility of their objections and inform the platform about the urgency of flagged content removal.
In addition, some participants desired an ability to provide feedback on their flagging experience (N = 6). That is, they wanted to voice their opinions not only about inappropriate posts on the platform but also about the flag submission process and its results. P249 elaborated on how the platform must listen to users’ opinions on the flagging mechanism:
@@
This reveals that participants want to contribute and express their perspectives on how the flagging mechanism should operate. In line with this, some participants recommended including a rebuttal or appeal option in the flagging process that would allow flaggers to contest the flag outcome decisions (N = 12). P708 specifically emphasized that an appeal process for flagging outcomes is crucial for both parties — the flagger and the author of the flagged post:
@@ 5.2.3.   Demanding Outcome Notifications with Timely Review
Many respondents (N = 284) expressed a strong preference for receiving notifications of flagging outcomes and explanations. Given that our survey questionnaire was designed to gauge perceptions of procedural fairness prior to decision-making for the flagged content, this finding highlights that participants closely associate the decisions regarding flagged posts with the overall fairness of the flagging mechanism. P312 emphasized the importance of transparency in decision-making, stating:
@@
These results contribute to our understanding that flagging outcomes impact perceived fairness not only for the authors whose posts were flagged for rule violations (Jhaver et al., 2019a) but also for the flaggers who reported the violations. In addition to participants who seek notification of flagging results, some (N = 46) expressed a desire for immediate post removal or suspension upon flagging. For instance, P304 suggested making posts invisible until a decision is reached on their removal.
@@
We observed that these participants emphasize the procedural fairness aspect of having their requests promptly addressed by the platform, associating ‘fairness’ with responsiveness to their concerns. In a similar vein, some participants (N = 31) expressed the need for an expedited review process to inform them of the results of their flagging promptly. P174 specifically noted the importance of timely reviews, recommending that:
Ensure follow-up or decision is made within 24 hours.
To sum up, users perceived fairness in the flagging process as being related to the platform’s immediate responsiveness and timely notification of the flagging results. From these opinions, we infer that users’ perceptions of fairness in flagging may depend on how seriously the platform takes their flagging requests and the time it takes to review the request.
@@ 5.2.4.   Expectations Regarding Review Procedures and Statistics
In addition to the flagging results, a significant portion of participants (N = 131) shared their need for more information on the flagging process, specifically about how the review process works. They argued that the flagger should be informed about the review criteria the platform adopts, who conducted the review, what steps are entailed in the review process, and how long they should expect to wait until the decision-making occurs. For instance, P500 proposed:
@@ 5.2.5.   Preventing Flagging Abuse and Protecting Flaggers
These insights align with findings from Zhang et al. (2023), highlighting that some users wish to have direct communication to effectively mitigate rule violations rather than submitting a flag, thereby achieving the educational purpose of rule enforcement.
@@ 6. Discussion
We began this study with the goal of examining how users engage with flagging mechanisms and how these mechanisms could be designed to enhance fairness perceptions among flaggers. Our mixed-methods analysis of responses from a large-scale survey suggests that many users recognize flagging as a critical tool to express their requests for removing inappropriate content online. Perhaps surprisingly, users’ engagement with flimsy flag icons, now ubiquitously available across most platforms (Crawford and Gillespie, 2016; Zhang et al., 2023), although often deeply embedded in action menus, triggers a range of sociopolitical concerns regarding platforms’ responsibilities, freedom of speech, algorithmic evaluation, safety against online harms, and privacy.
While flags offer a means for regular users to participate in the complex and opaque platform governance processes (Suzor, 2019; Ma et al., 2023a), users are keenly aware that flagging outcomes are largely determined by governance policies, mechanisms, and moderators instituted by platforms. This reflects the power and information asymmetry between users and platforms observed in prior research (Are, 2023) and emphasizes how platforms conceal, obscure, and strategically disclose details about their governance practices (Cotter, 2023).
From the users’ perspective, each aspect of flagging mechanisms — the procedural elements available during flagging, flag review criteria, flagging outcomes, and how they are communicated — require greater consistency, transparency, and support for user expression. Below, we discuss our specific design insights about how to achieve these goals, elaborate upon our conceptual contributions regarding flagging as a content moderation mechanism, and suggest promising avenues for further research.
@@ 6.1.1.   Incorporating a Text Box in Flagging Mechanisms
Our analysis reveals that flaggers’ fairness perceptions, especially regarding having a voice in the content moderation process, improve when they have an opportunity to express in detail their objections to the post during the flagging process. Specifically, our quantitative analysis demonstrates that the availability of a text box, which lets users articulate their thoughts in their own words within the flagging mechanism, significantly enhances their sense of being heard (Sec. 5.1.3). Contrary to our hypothesis H3-1 (Sec. 3.3), adding this text box to flag designs does not come at the cost of reducing users’ perceived consistency (Sec. 5.1.3). However, we also found that a text box increases the cognitive load on users when flagging a post (Sec. 5.1.5), which may reduce the mechanism’s overall usability. Although we did not find a direct association between this feature and users’ intention to flag again, prior research indicates that mental fatigue during the flagging process can lead to negative user experiences (Gillespie, 2018b). Therefore, platforms should carefully consider adopting this feature since it may limit user participation. One way for flag mechanisms to reduce unwarranted cognitive load could be to make user input in this text box optional and clarify that it is to be used only if flaggers want to provide additional necessary context.

### v2

@@ 5.1.4.   Moderator Type
To test Hypotheses 4, we conducted t-tests to examine how different moderator types affect user perceptions. Our analyses found that different moderator identities do not significantly impact perceived consistency, transparency, and voice. Therefore, we reject the hypotheses H4-1, H4-2, and H4-3.
5.1.5.   Additional Analyses: Interaction Effects and Usability
@@
Interaction Effects. In addition to testing our proposed hypotheses, we explored the interaction effects among the four flagging components on perceived consistency, transparency, and voice. Table 10 in Appendix C shows the results of this analysis. We estimated the statistical significance of these results following Bonferroni correction ($\alpha$05/11) and concluded that the interaction between classification schemes and the availability of a text box significantly affected perceived voice (F(2, 2936) = 8.29, p ¡ .001). Specifically, for participants in the condition without a text box, providing either a simple (MD = 0.34, SE = 0.10, p ¡ .001) or detailed (MD = 0.35, SE = 0.10, p ¡ .001) classification scheme elicited significantly higher perceptions of voice than providing no classification scheme (Table 11, Appendix C). All other interactions did not significantly impact any of the three aspects of procedural fairness.
Usability. We examined how variations in each component of flagging mechanisms differentially impacted the cognitive burden on our participants, thereby potentially influencing flag usability. Our analysis (Table 13, Appendix C) shows that participants in the condition with a text box (M = 3.86, SD = 1.73, EM = .05) experienced a significantly higher cognitive burden compared to those without such a box (M = 3.66, SD = 1.80, EM = .05). No other effects approached statistical significance. We also found that variations in how each flagging component is presented (or not presented) did not significantly affect participants’ self-reported likelihood of using flagging mechanism (Table 13, Appendix C).
@@ 5.2. Qualitative Findings
Our inductive analysis surfaced five themes through our combining and distilling of codes that together provide a nuanced understanding of user perspectives on enhancing fairness in flagging mechanisms, thus addressing RQ 4. Table 7 summarizes these themes and their frequency.
@@
Demanding outcome notifications with timely review 361 (29%)
@@ 5.2.1.   Desired Attributes of Flag Reviewers
Participants frequently expressed their concerns about flag reviewers’ biases and detailed their preferences regarding whether they prefer a human or AI moderator and their reasons for such preferences. Some participants (N = 61) mentioned that they prefer a human reviewer who can understand the context and nuance of the flagged post. Further, a few (N = 9) expressed their interest in communicating with moderators about the flagging process and the results. This suggests that users need thorough explanations for flagging outcomes through detailed conversations, revealing a preference for human interaction. For instance, participant P312 suggested:
@@
Many respondents offered suggestions to reduce biases in decision-making. For example, 17 participants suggested that platforms implement collaborative decision-making by involving multiple moderators for flagged posts. Some participants specified their suggestions for a certain combination or number of moderators (often three) or a mix of human and bot moderators. They believed that such cooperative decision-making in the flagging process would increase fairness and reduce biases. For example, P405 wrote:
@@
Further, to reduce a biased moderation review process, 27 participants emphasized the desired qualifications and characteristics of individual moderators. This includes moderators’ expertise in the subject matter of reviewed content, their understanding of ethical standards regarding rule violations, and their political neutrality.
@@ 5.2.2.   Needing Support for Greater Expressivity
Respondents often noted that the currently available flagging affordances do not allow them to fully communicate their objections to the flagged posts, and they proposed many constructive suggestions as to how flagging systems could better support expressivity. Among these, the most frequently suggested feature referenced the classification scheme, such as adopting a broader classification scheme (N = 45) with a wider range of rule violation categories for selections and the inclusion of multiple choices for flagging reasons (N = 90), which would enable flaggers to select several rule violation categories within the classification scheme. For example, Participant P158 advocated for allowing the selection of multiple categories, anticipating that it would induce changes at the flag review stage:
@@
Such responses reflect participants’ desire to classify and elaborate on the types of rule violations they encounter when making their reporting decisions. Other suggestions on how to achieve this included an ability to rate the severity of posts (N = 15) or highlight specific sections of posts that violate rules (N = 14). For instance, P708 noted:
@@
Further, some participants advocated for adopting measures that enable flaggers to corroborate a post’s toxicity by sharing their perspectives and referencing previously flagged posts. This finding suggests that incorporating multiple perspectives, not only from flag reviewers but also among flaggers themselves, could contribute to designing a fairer flagging mechanism. As specific methods, respondents expressed interest in discussing toxic posts in a forum (N = 21), where they could share their thoughts on the regulation of such posts. Another approach to ensuring accurate judgment of the post’s toxicity involved reviewing similar previously flagged posts and their post-review outcomes before submitting a flag. Participants believed that implementing such features would enhance the credibility of their objections and provide the platform with clear insights into the urgency of removing flagged content.
Besides communicating their objections to individual posts, some participants desired an ability to provide feedback on their flagging experience (N = 6). That is, they wanted to voice their opinions not only about inappropriate posts on the platform but also about the flag submission process and its results. P249 elaborated on how the platform must listen to users’ opinions on the flagging mechanism:
@@
This shows that participants want to contribute and express their perspectives on how the flagging mechanism should operate. In line with this, some participants recommended including a rebuttal or appeal option in the flagging process that would allow flaggers to contest the flag outcome decisions (N = 12). P708 specifically emphasized that an appeal process for flagging outcomes is crucial for both parties — the flagger and the author of the flagged post:
@@ 5.2.3.   Demanding Outcome Notifications with Timely Review
Many respondents (N = 284) expressed a strong preference for receiving notifications of flagging outcomes and explanations for those outcomes. Given that our survey questionnaire was designed to gauge perceptions of procedural fairness prior to decision-making for the flagged content, this finding highlights that participants closely associate the decisions regarding flagged posts with the overall fairness of the flagging mechanism. P312 emphasized the importance of transparency in decision-making, stating:
@@
Some respondents expressed a strong desire to receive a timely reply to their flags, which relates to the ‘voice’ aspect of procedural fairness. This suggests that flagging mechanisms should not only allow users to express their concerns during flagging but also ensure that these concerns are promptly considered and addressed in a responsive manner. This desire was especially reflected in participant suggestions for an expedited review process to inform flaggers of flag outcomes (N = 31). For instance, P174 noted the importance of timely reviews, recommending that:
Ensure follow-up or decision is made within 24 hours.
Additionally, some respondents (N = 46) called for immediate actions, such as post removal or account suspension upon flagging. For instance, P304 suggested making flagged posts invisible until a decision is reached about their removal:
@@
To sum up, users perceived fairness in the flagging process as being related to the platform’s immediate responsiveness and timely notification of the flagging results. We infer from these suggestions that users’ perceptions of fairness in flagging may depend on how seriously the platform takes their flagging requests, how clearly they explain the decision outcome, and the time it takes to review the requests.
@@ 5.2.4.   Expectations Regarding Review Procedures and Statistics
Respondents often desired to learn more about platforms’ processing of flags at the system-level. In addition to the flagging results, a significant portion of participants (N = 131) shared their need for more information on the flagging process, specifically about how the review process works. They argued that the flagger should be informed about the review criteria the platform adopts, who conducted the review, what steps are entailed in the review process, and how long they should expect to wait until the decision-making occurs. For instance, P500 proposed:
@@ 6. Discussion
We began this study with the goal of examining how flagging mechanisms should be designed to enhance fairness perceptions among flaggers. Our statistical analysis of responses from a large-scale survey experiment shows that including posting guidelines and a text box for feedback within flag implementations helps enhance users’ fairness perceptions, whereas offering classification schemes or providing information about whether the flag reviewer is a human or a bot does not significantly influence users attitudes. Our qualitative analysis of open-ended responses shows that users feel concerned about reviewers’ biases, desire flagging systems to support greater expressivity, demand timely notifications and explanations of flag outcomes, wish to see information about flag processing at the system-level, and expect platforms to prevent flag abuse.
These results contribute empirically informed guidance on how social media platforms should design different components of their flagging interfaces and how these design choices could impact users’ attitudes toward flagging. We document the key information and security needs of flag submitters and offer insights for how platforms could address them. We show that users’ engagement with flags triggers a range of sociopolitical concerns regarding platforms’ responsibilities, freedom of speech, algorithmic evaluation, safety against online harms, and privacy. We also provide a methodological framework that others can adopt to evaluate the effectiveness of new components and affordances that seek to address users’ reporting needs.
Prior empirical research on enacting fairness in content moderation (Myers West, 2018; Jhaver et al., 2019a, c; Ma et al., 2023b; Vaccaro et al., 2020) has largely focused on how moderation decisions are enacted and communicated to moderated users. Our work contributes to this research by examining procedural fairness during flag submissions. We show that from the flaggers’ perspective, each aspect of flagging mechanisms — the procedural elements available during flagging, flag review criteria, flagging outcomes, and how they are communicated — require greater consistency, transparency, and support for user expression. Below, we discuss our specific design insights about how to achieve these goals, elaborate upon our conceptual contributions regarding flagging as a content moderation mechanism, and suggest promising avenues for further research.
@@ 6.1.1.   Incorporating a Text Box in Flagging Mechanisms
Our analysis reveals that flaggers’ fairness perceptions, especially regarding having a voice in the content moderation process, improve when they have an opportunity to express in detail their objections to the post during the flagging process. Specifically, our quantitative analysis demonstrates that the availability of a text box, which lets users articulate their thoughts in their own words within the flagging mechanism, significantly enhances their sense of being heard (Sec. 5.1.3). Contrary to our hypothesis H3-1 (Sec. 3.3), adding this text box to flag designs does not come at the cost of reducing users’ perceived consistency (Sec. 5.1.3). Besides procedural fairness, a text box could also enhance outcome fairness by providing flag reviewers additional context to fairly evaluate flagged posts. Flag outcomes that are accompanied by explanations regarding decision-making (Jhaver et al., 2019a, c) and how flaggers’ voice was considered could improve flaggers’ and flagged users’ perceptions of both procedural and outcome fairness.
We found that a text box increases the cognitive load on users when flagging a post (Sec. 5.1.5), which may reduce the mechanism’s overall usability. Although we did not find a direct association between this feature and users’ intention to flag again, prior research indicates that mental fatigue during the flagging process can lead to negative user experiences (Gillespie, 2018b). Therefore, platforms should carefully consider adopting this feature since it may limit user participation. One way for flag mechanisms to reduce unwarranted cognitive load could be to make user input in this text box optional and clarify that it is to be used only if flaggers want to provide additional necessary context.